OpenStudy (anonymous):

Biscuits are sold 36 per case, and sausage patties are sold 42 per case. Find the least common multiple. Then determine how many cases of each Billy should buy to ensure no extras when all biscuits are sold as sausage biscuits.

OpenStudy (anonymous):

do you know how to find the least common multiple of two numbers? it helps if they are in factored form

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok, then since \(36=2^2\times 3^2\) and \(42=2\times 3\times 7\) the least common multiple is \(2^2\times 3^2\times 7=252\)

OpenStudy (anonymous):

So he should buy 252 of each?

OpenStudy (anonymous):

careful here.

OpenStudy (anonymous):

the least common multiple of 36 and 42 is 252, but there are 36 biscuits per case, so he should be 7 cases of biscuits, because \(7\times 36=252\)

OpenStudy (anonymous):

and similarly he should be 6 cases of sausages, because \(42\times 6=252\)

OpenStudy (anonymous):

Ok thank you SOOO much!!!!!!!!!!!!!!

OpenStudy (anonymous):

yw